# A Survey on Large Language Model Acceleration based on KV Cache Management

### v2

@@ 4.4.2 Mixed-precision quantization
Unlike fixed-precision quantization, where all Keys or Values are quantized to the same bit-width (e.g., 4-bit or 8-bit), mixed-precision quantization assigns higher or full precision to Keys and Values of critical tokens and parts while using lower precision for less critical parts. The summary of KV mixed-precision quantization is listed in Tab. V. KVQuant [84] proposes several strategies to quantize Keys and Values smoothly based on observations. Firstly, KVQuant observes that the key values exhibit outliers in specific channels prior to applying Rotary Positional Embedding (RoPE). However, after applying RoPE, the magnitudes of these outlier channels become less consistent, creating a unique challenge for low-precision quantization. Thus, KVQuant [84] proposes to quantize the Keys per channel before applying the RoPE operations and to quantize the Values per token. Secondly, KVQuant [84] observes that KV cache activations contain outliers that skew the quantization range. To address this, they isolate outliers per vector (e.g., per-channel or per-token), store them in a sparse format, and quantize the remaining values to a narrower range. Thirdly, LLMs disproportionately allocate high attention scores to the first token (i.e., attention sink), and quantizing the first token will damage the performance of LLMs. Thus, KVQuant [84] retains the first token in full precision (FP16) while quantizing the rest of the sequence, which is also used by IntactKV [85] and SKVQ [86]. Similar to KVQuant [84], KIVI [87] quantizes the Key cache per-channel, as certain channels exhibit large outliers, and the Value cache per-token, as there are no significant outlier patterns in the Value cache. Additionally, KIVI [87] retains the most recent Keys and Values in full precision, while quantizing older KVs. This approach is based on the observation that the most recent KVs are critical for generating subsequent tokens.
@@
QAQ [148] proposes a quality adaptive quantization approach to dynamically determine the suitable quantization bit for each token, based on its importance and sensitivity, while handling outliers and exceptions to maintain model performance. SKVQ [86] introduces the clipped dynamic quantization with channel reorder. First, SKVQ [86] uses a transformation-invariant permutation to group similar channels based on their statistical characteristics and applies clipped dynamic quantization to mitigate the outlier problem. Second, SKVQ [86] maintains high precision for the initial tokens and the most recent tokens while quantizing older tokens. Consequently, SKVQ [86] effectively reduces quantization errors and improves the accuracy of the quantized model. CacheGen [149] and AsymKV [167] use layer-wise asymmetric quantization, assigning higher-bit precision to key matrices in sensitive early layers and lower-bit precision to less sensitive layers, balancing memory efficiency and performance. Particularly, CacheGen [149] also exploits token-wise locality by encoding deltas (differences) between KV tensors of nearby tokens instead of raw values. Atom [150] identifies and separates outlier channels, reordering the matrix to group these outlier channels at the end, thereby ensuring regular memory access patterns for improved hardware utilization. Then, Atom [150] quantizes outliers with higher precision, while normal channels are quantized to INT4 for maximum efficiency. In particular, Atom [150] applies fine-grained group quantization by dividing matrices into smaller subgroups (e.g., 128 elements per group) and performing quantization independently within each group.
@@ 4.4.3 Outlier Redistribution
Specifically, MassiveActivation [72] highlights the phenomenon of massive activations in large language models (LLMs), where a small subset of activations is exponentially larger than the rest. To address this, MassiveActivation [72] proposes appending a virtual token to the inputs, allowing LLMs to encapsulate the massive outliers within these learned keys and values for each head. Then, we introduce the equivalent transformation function-based approaches. Firstly, QuaRot [73], Qserve [74], and Q-INT4 [75] redistributes outlier values across all channels by Hadamard rotation, successfully lowering the maximum value of outlier tokens. The Hadamard rotation of activations can be incorporated into the preceding linear layer, thereby redistributing the outliers of Keys and Values into the parameters. Despite this improvement, outlier tokens still exhibit magnitudes hundreds of times greater than normal tokens, causing notable performance issues when using shared quantization scales across tokens [93]. Expanding on this idea, SpinQuant [76] proposes training an orthogonal matrix instead of relying on a random Hadamard matrix to achieve better performance. Similarly, DuQuant [77] employs channel permutation to evenly distribute outliers across blocks and utilizes block rotation to further smooth outliers.
@@
Instead of using handcrafted transformations [77, 79, 78] to shift the quantization difficulty from activations to weights, AffineQuant [80] uses an affine transformation matrix that combines both scaling and rotation transformations. This allows it to optimize weight distributions more effectively, aligning them better with the quantization function and reducing quantization errors. The affine transformation matrix provides richer flexibility compared to SmoothQuant’s scalar-based scaling, enabling finer adjustments to the weight and activation distributions. Based on AffineQuant [80], FlatQuant [81] introduces a fast and learnable affine transformations to enhance the flatness of weights and activations, which decomposes transformations into smaller matrices to reduce memory and computational costs. Similarly, AWQ [82] and OmniQuant [83] proposes differentiable and learnable equivalent transformations, which optimize the equivalent parameters (e.g., channel-wise scaling and shifting) in an end-to-end manner using gradient descent.
@@ 4.4.4 Summary and Future Directions
KV cache quantization is a crucial technique for reducing memory and computational overhead in large language models (LLMs) during autoregressive decoding. While significant progress has been made, this field remains dynamic and rapidly evolving, with several promising directions for future research. Firstly, one promising avenue is the development of real-time adaptive quantization methods. These techniques could dynamically adjust quantization levels during inference based on real-time metrics such as token importance, outlier presence, or sequence length. Such an approach could significantly enhance efficiency while maintaining performance, especially for processing long sequences with varying levels of complexity. Secondly, another important direction is extending KV cache quantization to multi-modal and multi-task models. Multi-modal models, which process inputs from diverse domains such as text, vision, and audio, and multi-task scenarios often exhibit highly diverse attention patterns and memory demands. This necessitates the design of more advanced and tailored quantization strategies to balance efficiency and accuracy in these increasingly complex settings.
@@ 4.5.1 Singular Value Decomposition
Firstly, ECKVH [66], EigenAttention [67], and ZDC [68] shows that KV caches have a low-rank property, where a small number of top singular values retain most of the information. Using Singular Value Decomposition (SVD), the method compresses KV caches by grouping heads, applying SVD, and retaining top singular values, effectively reducing the number of KV heads with minimal error. Also, ZDC [68] uses an adaptive hybrid compression ratio mechanism to assign higher compression to unimportant tokens in shallower layers while preserving more important tokens in deeper layers, leveraging the similarity of token characteristics in adjacent layers. Secondly, rather than decomposing KV pairs, LoRC [69] employs a low-rank approximation of KV weight matrices and adopts a progressive compression strategy to efficiently compress KV caches without requiring model retraining. Specifically, LoRC [69] uses SVD to compress the Keys and Values parameter matrices (i.e., $\mathbf{W}^{k}_{i}$ and $\mathbf{W}^{v}_{i}$) as $\mathbf{W}^{k}_{i}=\mathbf{U}^{k}_{i}\mathbf{\Sigma}^{k}_{i}{\mathbf{P}^{k}_{i}}^{\top}$ and $\mathbf{W}^{v}_{i}=\mathbf{U}^{v}_{i}\mathbf{\Sigma}^{v}_{i}{\mathbf{P}^{v}_{i}}^{\top}$. Also, compression is applied conservatively in shallower layers to minimize error amplification and more aggressively in deeper layers. Then, instead of storing $\mathbf{K}^{i}=\mathbf{X}_{i}\mathbf{W}^{k}_{i}$ and $\mathbf{V}^{i}=\mathbf{X}_{i}\mathbf{W}^{v}_{i}$, it only stores $\mathbf{\hat{K}}^{i}=\mathbf{X}_{i}\mathbf{U}^{k}_{i}$ and $\mathbf{\hat{V}}^{i}=\mathbf{X}_{i}\mathbf{U}^{v}_{i}$, along with $\mathbf{\Sigma}^{k}_{i}{\mathbf{P}^{k}_{i}}^{\top}$ and $\mathbf{\Sigma}^{v}_{i}{\mathbf{P}^{v}_{i}}^{\top}$. Also, ShadowKV [70] performs SVD decomposition directly on pre-RoPE keys to reduce the dimensionality of the key representations. Palu [71] applies SVD to compress both Keys and Values.
@@ 4.5.2 Tensor Decomposition
where $\mathcal{T}_{(k)}$ represents the local tensor of size $d_{k-1}\times i_{k}\times j_{k}\times d_{k}$, with $\prod_{k=1}^{n}i_{k}=I$ and $\prod_{k=1}^{n}j_{k}=J$. Here, $n$ denotes the number of local tensors, collectively referred to as the decomposed tensors. As shown in Eaquation (12), MPO-based tensor decomposition is well-suited for KV cache compression as it reduces the memory footprint by factorizing large key and value matrices into smaller local tensors, enabling efficient storage while preserving essential information. This approach minimizes redundancy and maintains the structural integrity required for accurate attention computations. DecoQuant [65] combines quantization with low-rank decomposition to effectively reduce quantization errors. Specifically, DecoQuant [65] leverages the Matrix Product Operator (MPO) to decompose matrices into smaller local tensors. The larger tensors, which contain most of the parameters, are quantized to low-bit precision, while the smaller tensors retain high precision to minimize overall quantization error.
@@ 5 Model-level Optimization
In model-level optimization, new architectures or mechanisms are designed for transformers to allow more efficient reuse of KV cache. Typically, these methods require retraining or fine-tuning of the model to come into operation. Nevertheless, efficient transformation pipelines have also been proposed to allow for a fast deployment to new architectures. According to where and how the refinement was made to the models, we separate related works to the grouping and sharing mechanisms within or cross layers (Sec. 5.1), implementing architecture modification or augmentation (Sec. 5.2), and incorporating non-transformer architectures for optimization (Sec. 5.3). The taxonomy of the model-level optimization is shown in Fig. 5.
@@ 5.1 Attention Grouping and Sharing
This section explores attention grouping and sharing methods as effective strategies for optimizing key-value (KV) management. We categorize the approaches into two distinct subtypes: intra-layer grouping (Sec. 5.1.1) that focuses on grouping query, key, and value heads within individual layers to reduce redundancy and improve efficiency, and cross-layer sharing 5.1.2 that shares key, value, or attention components across layers to improve information reuse and reduce KV cache requirements. The summary of attention grouping and sharing is listed in Tab. VII.
@@ 5.1.1 Intra-layer Grouping
Shazeer first introduced Multi-Query Attention (MQA) [195] that modified the traditional multi-head attention mechanism. In MQA, all attention heads in a transformer block share a single key and value. This simple strategy can greatly accelerate the decoding procedure. The experiments of the author show that MQA would gain much efficiency with only minor quality degradation incurring.
@@
There were several extensions based on GQA. AsymGQA [197] extends GQA by proposing an activation-informed merging strategy. Instead of grouping the heads by uniform clustering, AsymGQA dynamically determines the grouping of quries based on their activations similarities during training and constructs an asymmetric group results, which leads to better optimization and generalization. Weighted GQA [198] introduces additional trainable weights to each key and value head, which can be seamlessly integrated into existing GQA models. By tuning weights during training, it improves the performance of the model without additional inference overhead. QCQA [199] utilizes an evolutionary algorithm to identify the optimal query head groupings for GQA, which is guided by a computationally efficient fitness function that leverages the weight-sharing error and the KV cache to evaluate text generation quality and memory capacity. KDGQA [200] argues that many variances of GQA adopt a fixed grouping strategy, thus lacking dynamic adaptability to the evolving of key-value interactions during training. Their Dynamic Key-Driven GQA address these issues by allocating groups using key head norms adaptively during training, resulting in a flexible strategy to query head grouping and enhance the performance.
GQKVA [201] advances the grouping strategy and comes up with a generalized query, key and value grouping mechanism. It first introduces MKVA and GKVA, in which the key and value are grouped to share the same query. Based on this, GQKVA is proposed to separately group the query and key-value pairs. Typically, queries are partitioned into $g_{q}$ groups, and keys and values are partitioned into $g_{kv}$ groups, and each combination of query and key-value pairs would interact using dot product attention. This results in $g_{q}\times g_{kv}$ distinct outputs. It generalized different group strategy on query, key and value and preserves good computational efficiency and comparable performance as MHA.
@@ 5.1.2 Cross-layer Sharing
Brandon et al. introduce Cross Layer Attention (CLA) [186] that extends the ideas of GQA and MQA by sharing the key and value heads between adjacent layers, further reduce the redundancy in the KV cache. This achieves an additional 2$\times$ KV cache size reduction compared to MQA, significantly improving memory efficiency without altering computational complexity.
LCKV [187] proposes only to compute and cache the key and value for a small subset of layers, even only the top layer, then let queries in bottom layers pair the saved keys and values for inference. This method not only drastically improves the inference speed and reduces memory consumption but is also orthogonal to existing memory-saving techniques, enabling straightforward integration for further optimization. While such a mechanism makes next token computation depend on top layer keys and values of previous tokens, which contradict to the parallel training of transformers, LCKV introduces an approximate training methods to support parallel training.
@@
LISA (Lightweight Substitute for Attention) [190] makes a comprehensive analysis for the similarity of attention patterns across layers. Directly sharing attention weights across layers is ineffective because of the misalignment of the attention head and the sensitivity of shallow layers. LISA [190] addresses challenges by incorporating tiny feed-forward networks to align attention heads between layers and using low-rank matrices to approximate variations in layer-wise attention weights. This achieves a 6× compression of query and key parameters while maintaining high accuracy and perplexity.
Wu et al. [191] introduce a unified framework that systematically analyzes and optimizes the cross-layer Key-Value cache sharing mechanism. They consolidate several existing methods, explore novel variants within a cohesive structure, and make thorough evaluations of these methods. The study finds that 2 times reduction to KV cache size can outperform standard transformers in throughput without substantial accuracy loss, while further reduction requires alternative design with additional training costs. With the analysis results, they offer insight into the choice of appropriate KV sharing methods based on the specific requirement or constraints.

### v3

@@ 4.4.2 Mixed-precision quantization
Unlike fixed-precision quantization, where all Keys or Values are quantized to the same bit-width (e.g., 4-bit or 8-bit), mixed-precision quantization assigns higher or full precision to Keys and Values of critical tokens and parts while using lower precision for less critical parts. The summary of KV mixed-precision quantization is listed in Tab. V. KVQuant [87] proposes several strategies to quantize Keys and Values smoothly based on observations. Firstly, KVQuant observes that the key values exhibit outliers in specific channels prior to applying Rotary Positional Embedding (RoPE). However, after applying RoPE, the magnitudes of these outlier channels become less consistent, creating a unique challenge for low-precision quantization. Thus, KVQuant [87] proposes to quantize the Keys per channel before applying the RoPE operations and to quantize the Values per token. Secondly, KVQuant [87] observes that KV cache activations contain outliers that skew the quantization range. To address this, they isolate outliers per vector (e.g., per-channel or per-token), store them in a sparse format, and quantize the remaining values to a narrower range. Thirdly, LLMs disproportionately allocate high attention scores to the first token (i.e., attention sink), and quantizing the first token will damage the performance of LLMs. Thus, KVQuant [87] retains the first token in full precision (FP16) while quantizing the rest of the sequence, which is also used by IntactKV [88] and SKVQ [89]. Similar to KVQuant [87], KIVI [90] quantizes the Key cache per-channel, as certain channels exhibit large outliers, and the Value cache per-token, as there are no significant outlier patterns in the Value cache. Additionally, KIVI [90] retains the most recent Keys and Values in full precision, while quantizing older KVs. This approach is based on the observation that the most recent KVs are critical for generating subsequent tokens. Recent work [174] provides mathematical justification for prioritizing key precision over value precision through spectral norm analysis.
@@
KV-AdaQuant [177] observes the higher sensitivity of key matrices to quantization errors due to their larger norms, and therefore allocates more bits to key matrices and fewer bits to value matrices, optimizing both accuracy and memory efficiency. QAQ [156] proposes a quality adaptive quantization approach to dynamically determine the suitable quantization bit for each token, based on its importance and sensitivity, while handling outliers and exceptions to maintain model performance. SKVQ [89] introduces the clipped dynamic quantization with channel reorder. First, SKVQ [89] uses a transformation-invariant permutation to group similar channels based on their statistical characteristics and applies clipped dynamic quantization to mitigate the outlier problem. Second, SKVQ [89] maintains high precision for the initial tokens and the most recent tokens while quantizing older tokens. Consequently, SKVQ [89] effectively reduces quantization errors and improves the accuracy of the quantized model. CacheGen [157] and AsymKV [178] use layer-wise asymmetric quantization, assigning higher-bit precision to key matrices in sensitive early layers and lower-bit precision to less sensitive layers, balancing memory efficiency and performance. Particularly, CacheGen [157] also exploits token-wise locality by encoding deltas (differences) between KV tensors of nearby tokens instead of raw values. Atom [158] identifies and separates outlier channels, reordering the matrix to group these outlier channels at the end, thereby ensuring regular memory access patterns for improved hardware utilization. Then, Atom [158] quantizes outliers with higher precision, while normal channels are quantized to INT4 for maximum efficiency. In particular, Atom [158] applies fine-grained group quantization by dividing matrices into smaller subgroups (e.g., 128 elements per group) and performing quantization independently within each group.
@@ 4.4.3 Outlier Redistribution
Specifically, MassiveActivation [75] highlights the phenomenon of massive activations in large language models (LLMs), where a small subset of activations is exponentially larger than the rest. To address this, MassiveActivation [75] proposes appending a virtual token to the inputs, allowing LLMs to encapsulate the massive outliers within these learned keys and values for each head. On the other hand, to further address this issue, existing researchers [76, 77] introduce equivalent transformation function-based approaches. Equivalent transformation functions are mathematical transformations applied to activations that preserve the underlying information while redistributing or normalizing the values. By redistributing or scaling massive values, these approaches ensure that extreme outliers do not disproportionately affect downstream processes like compression or quantization. Firstly, QuaRot [76], Qserve [77], and Q-INT4 [78] redistributes outlier values across all channels by Hadamard rotation, successfully lowering the maximum value of outlier tokens. The Hadamard rotation of activations can be incorporated into the preceding linear layer, thereby effectively redistributing the outliers of Keys and Values into the parameters. Despite this improvement, outlier tokens still exhibit magnitudes hundreds of times greater than normal tokens, causing notable performance issues when using shared quantization scales across tokens [96]. Expanding on this idea, SpinQuant [79] proposes training an orthogonal matrix instead of relying on a random Hadamard matrix to achieve better performance. Similarly, DuQuant [80] employs channel permutation to evenly distribute outliers across blocks and utilizes block rotation to further smooth outliers.
@@
Instead of using handcrafted transformations [80, 82, 81] to shift the quantization difficulty from activations to weights, AffineQuant [83] uses an affine transformation matrix that combines both scaling and rotation transformations. This allows it to optimize weight distributions more effectively, aligning them better with the quantization function and reducing quantization errors. The affine transformation matrix provides richer flexibility compared to SmoothQuant’s scalar-based scaling, enabling finer adjustments to the weight and activation distributions. Based on AffineQuant [83], FlatQuant [84] introduces a fast and learnable affine transformation to enhance the flatness of weights and activations, which intelligently decomposes transformation into smaller matrices to reduce memory and computational costs. Similarly, AWQ [85] and OmniQuant [86] propose differentiable and learnable equivalent transformations, which optimize the equivalent parameters (e.g., channel-wise scaling and shifting) in an end-to-end manner using gradient descent.
@@ 4.4.4 Summary and Future Directions
KV cache quantization is a crucial technique for reducing memory and computational overhead in large language models (LLMs) during auto-regressive decoding. While significant progress has been made, this field remains dynamic and rapidly evolving, with several promising directions for future research. Firstly, one promising avenue is the development of real-time adaptive quantization methods. These techniques could dynamically adjust quantization levels during inference based on real-time metrics such as token importance, outlier presence, or sequence length. Such an approach could significantly enhance efficiency while maintaining performance, especially for processing long sequences with varying levels of complexity. Secondly, another important direction is extending KV cache quantization to multi-modal and multi-task models. Multi-modal models, which process inputs from diverse domains such as text, vision, and audio, and multi-task scenarios often exhibit highly diverse attention patterns and memory demands. This necessitates the design of more advanced and tailored quantization strategies to balance efficiency and accuracy in these increasingly complex settings.
@@ 4.5.1 Singular Value Decomposition
Recent research has demonstrated that KV caches in large language models exhibit strong low-rank properties, where a small number of top singular values retain most of the information. Building upon this discovery, numerous low-rank approximation techniques have been proposed for KV cache optimization. These methods can be broadly categorized into two approaches: those that directly decompose the KV cache and those that apply low-rank approximations to the KV weight matrices.
Among the methods that directly decompose KV caches, ECKVH [67], EigenAttention [68], Q-Filters [73], and ZDC [69] all leverage Singular Value Decomposition (SVD) techniques but with distinct implementations. ECKVH [67] compresses KV caches by grouping attention heads, applying SVD, and retaining top singular values, which effectively reduces the number of KV heads while minimizing error. Similarly, EigenAttention [68] strategically employs SVD to effectively approximate keys, queries, and values with low-rank basis vectors, thereby reducing the dimensionality of KV matrices. Q-Filters [73], meanwhile, innovatively offers a training-free KV cache compression method that utilizes SVD-based key projections for efficient and accurate attention score approximations.
Besides, Loki [74] introduces an alternative approach that first computes approximate attention scores in a reduced dimensional space to efficiently rank and select top-k keys. It then calculates exact attention scores using only the selected keys in the transformed space, thereby reducing computational and memory requirements while maintaining performance. Furthermore, ZDC [69] implements an adaptive hybrid compression ratio mechanism that assigns higher compression to less important tokens in shallower layers while preserving more important tokens in deeper layers, effectively leveraging the similarity of token characteristics in adjacent layers.
Taking a fundamentally different approach, LoRC [70] employs low-rank approximations of KV weight matrices rather than decomposing the KV pairs themselves. Specifically, LoRC uses SVD to compress the Keys and Values parameter matrices (i.e., $\mathbf{W}^{k}_{i}$ and $\mathbf{W}^{v}_{i}$), decomposing them as $\mathbf{W}^{k}_{i}=\mathbf{U}^{k}_{i}\mathbf{\Sigma}^{k}_{i}{\mathbf{P}^{k}_{i}}^{\top}$ and $\mathbf{W}^{v}_{i}=\mathbf{U}^{v}_{i}\mathbf{\Sigma}^{v}_{i}{\mathbf{P}^{v}_{i}}^{\top}$. Additionally, LoRC adopts a progressive compression strategy, applying compression conservatively in shallower layers to minimize error amplification while compressing more aggressively in deeper layers. Instead of storing complete $\mathbf{K}^{i}=\mathbf{X}_{i}\mathbf{W}^{k}_{i}$ and $\mathbf{V}^{i}=\mathbf{X}_{i}\mathbf{W}^{v}_{i}$ matrices, it only stores $\mathbf{\hat{K}}^{i}=\mathbf{X}_{i}\mathbf{U}^{k}_{i}$ and $\mathbf{\hat{V}}^{i}=\mathbf{X}_{i}\mathbf{U}^{v}_{i}$, along with $\mathbf{\Sigma}^{k}_{i}{\mathbf{P}^{k}_{i}}^{\top}$ and $\mathbf{\Sigma}^{v}_{i}{\mathbf{P}^{v}_{i}}^{\top}$, achieving efficient KV cache compression without requiring model retraining. Palu [72] follows a comparable approach by applying SVD to compress KV weight matrices simultaneously, while ShadowKV [71] takes a different angle by performing SVD decomposition directly on pre-RoPE keys to reduce the dimensionality of key representations.
@@ 4.5.2 Tensor Decomposition
where $\mathcal{T}_{(k)}$ represents the local tensor of size $d_{k-1}\times i_{k}\times j_{k}\times d_{k}$, with $\prod_{k=1}^{n}i_{k}=I$ and $\prod_{k=1}^{n}j_{k}=J$. Here, $n$ denotes the number of local tensors, collectively referred to as the decomposed tensors. As shown in Equation 12, MPO-based tensor decomposition is well-suited for KV cache compression as it reduces the memory footprint by factorizing large key and value matrices into smaller local tensors, enabling efficient storage while preserving essential information. This approach minimizes redundancy and maintains the structural integrity required for accurate attention computations. DecoQuant [66] combines quantization with low-rank decomposition to effectively reduce quantization errors. Specifically, DecoQuant [66] leverages the Matrix Product Operator (MPO) to decompose matrices into smaller local tensors. The larger tensors, which contain most of the parameters, are quantized to low-bit precision, while the smaller tensors retain high precision to minimize overall quantization error.
@@ 5 Model-level Optimization
In model-level optimization, new architectures or mechanisms are designed for transformers to allow more efficient reuse of KV cache. Typically, these methods require retraining or fine-tuning of the model to come into operation. Nevertheless, efficient transformation pipelines have also been proposed to allow for a fast deployment to new architectures. We categorize related works according to their grouping and sharing mechanisms, either within layers or across layers (Sec. 5.1), implementing architecture modification or augmentation (Sec. 5.2), and incorporating non-transformer architectures for optimization (Sec. 5.3). The taxonomy of the model-level optimization is shown in Fig. 7.
@@ 5.1 Attention Grouping and Sharing
This section explores attention grouping and sharing methods as effective strategies for optimizing key-value (KV) management. We categorize the approaches into two distinct subtypes: intra-layer grouping (Sec. 5.1.1) that focuses on grouping query, key, and value heads within individual layers to reduce redundancy and improve efficiency, and cross-layer sharing (Sec. 5.1.2) that shares key, value, or attention components across layers to improve information reuse and reduce KV cache requirements. The summary of attention grouping and sharing is listed in Tab. VII.
@@ 5.1.1 Intra-layer Grouping
Figure 8: Different QKV techniques.
As shown in Fig. 8, Shazeer first introduced Multi-Query Attention (MQA) [206] that modified the traditional multi-head attention mechanism. In MQA, all attention heads in a transformer block share a single key and value. This simple strategy can greatly accelerate the decoding procedure. The experiments of the author show that MQA would gain much efficiency with only minor quality degradation occurring.
@@ 5.1.2 Cross-layer Sharing
Brandon et al. introduce Cross Layer Attention (CLA) [197] that extends the ideas of GQA and MQA by sharing the key and value heads between adjacent layers, further reducing the redundancy in the KV cache. This achieves an additional 2$\times$ KV cache size reduction compared to MQA, significantly improving memory efficiency without altering computational complexity.
LCKV [198] proposes only to compute and cache the key and value for a small subset of layers, even only the top layer, then let queries in bottom layers pair the saved keys and values for inference. This method not only drastically improves the inference speed and reduces memory consumption but is also orthogonal to existing memory-saving techniques, enabling straightforward integration for further optimization. While such a mechanism makes next token computation depend on top layer keys and values of previous tokens, which contradicts the parallel training of transformers, LCKV introduces an approximate training method to support parallel training.
@@
LISA (Lightweight Substitute for Attention) [201] makes a comprehensive analysis of the similarity of attention patterns across layers. Directly sharing attention weights across layers is ineffective because of the misalignment of the attention head and the sensitivity of shallow layers. LISA [201] addresses challenges by incorporating tiny feed-forward networks to align attention heads between layers and using low-rank matrices to approximate variations in layer-wise attention weights. This remarkably achieves a 6× compression of query and key parameters while maintaining high accuracy and perplexity.
Wu et al. [202] introduce a unified framework that systematically analyzes and optimizes the cross-layer Key-Value cache sharing mechanism. They consolidate several existing methods, explore novel variants within a cohesive structure, and make thorough evaluations of these methods. The study finds that two times reduction to KV cache size can outperform standard transformers in throughput without substantial accuracy loss, while further reduction requires alternative design with additional training costs. With the analysis results, they offer insight into the choice of appropriate KV sharing methods based on the specific requirements or constraints.
DHA (Decoupled Head Attention) [204] addresses redundancy in MHA and adaptively configures shared groups for key and value heads across layers, reducing KV cache requirements. Observing that clustering and fusing similar heads can reduce KV cache size without significant performance reduction, DHA designs a search, fusion, and continued pre-training framework that can progressively transform MHA checkpoints into DHA models through linear fusion of head parameters, preserving the pre-trained knowledge with a small pre-training budget.
5.1.3 Summary and Future Directions
This section explores innovative strategies for optimizing memory and computational efficiency through intra-layer grouping and cross-layer sharing mechanisms, while identifying key challenges and future directions. Maintaining performance for precision-sensitive tasks, ensuring scalability across diverse model architectures, and addressing the under-explored dynamics of attention across time and layers remain critical areas for improvement. Current approaches, like DHA [204] and LISA [201], often struggle to generalize to emerging or non-standard architectures, while static grouping and sharing strategies fail to capture temporal and contextual attention variations. Future research should focus on developing universal, adaptable frameworks that require minimal retraining, integrating optimization techniques like quantization and pruning, and leveraging dynamic, runtime adjustments to better capture task-specific requirements. Additionally, understanding the downstream impacts on fine-tuning and transfer learning is essential for practical real-world applicability.